# Learning to be Indifferent in Complex Decisions:

### v1

@@ 3.1 Example: Multiplicity of SVE
    Figure 4: Stable strict pure VE at $(1.0,0.0)$; $\beta=50$
    Figure 5: Stable strict pure VE at $(0.0,0.5)$; $\beta=50$
@@ 3.3 Example: Unique Pure SVE
Figure 6: Stable Unique Mixed VE at $(2.577,2.577)$; $\beta=50$
Figure 7: Stable Unique Strict Pure VE at $(1.5,0)$; $\beta=50$
It is important to note that if the decision-maker possesses the finest partition available (i.e., she can distinguish each alternative independently), then multiple equilibria and mixed equilibria are impossible in a decision tree with generic payoffs in our learning model. Thus, the clustering of alternatives has non-trivial consequences on the dynamics of an otherwise standard reinforcement learning model. Finally, the examples suggest that an asymptotically stable equilibrium of the CPAL model always exists, and that strict pure equilibria are asymptotically stable whenever they are present. In the next section, we formalize these insights and provide general results on the convergence of CPAL dynamics when the DM operates with at most two similarity classes. Later in the paper, we analyze the general case with an arbitrary number of similarity classes. [^19] Supplementary examples showing the workings of the CPAL model with more than two similarity classes can be found in the Online Appendix.
@@ 4 Decision Trees with Two Similarity Classes
There exists a finite $\hat{\beta}$, such that $\forall$ $\beta>\hat{\beta}$, a decision tree $\mathcal{T^{\prime}}$ with generic payoffs where an agent chooses among alternatives in at most two similarity classes admits a smooth valuation equilibrium that is locally asymptotically stable in the CPAL dynamics. Additionally, if the equilibrium is unique, it is globally asymptotically stable.
@@ 5 Decision Trees with More than Two Similarity Classes
We extend our analysis to decision problems with generic payoffs where Alice has an arbitrary number of similarity classes available to her. Specifically, we let $|\mathcal{S}|=n$ where $n\in\mathbb{N}:n>2$, although we still maintain that $n$ is finite and that at any node $\omega$, $1\leq|\mathcal{S}_{\omega}|\leq n$. Once again, we imagine a scenario where at the parent node $r^{\prime}$, nature selects a node $\omega\in\Omega$ with probability $p(\omega)\geq 0$ such that $\sum_{\omega\in\Omega}p(\omega)=1$ where $\Omega={\mathcal{P(S)}}\setminus\{\emptyset\}$. For Theorem 3, we additionally assume that the set of nodes $\{\omega\in\Omega:|\mathcal{S}_{\omega}|=1\}$ is included in the support of the probability mass function $p$. That is, for each similarity class $s\in\mathcal{S}$, the trivial unary choice node featuring it is drawn by nature with a strictly positive probability. We recall that Theorem 1 guarantees the existence of a steady-state of the CPAL model and Lemma 2.1 tells us that a steady-state may arise in an arbitrarily small neighborhood of either a strict pure VE, a partially-mixed VE or a fully-mixed VE for a sufficiently large sensitivity parameter. We use the notation introduced in Section 2.3.
@@ 5.2 When Trivial Choice Payoffs are Large
Proof in the Online Appendix. ∎
@@ 5.3 When Trivial Choice Payoffs are Small
To illustrate the dramatic impact of assuming large trivial choice payoffs, we now consider the opposite scenario where these payoffs are small, leading to markedly different properties for the steady-states of the CPAL dynamics. Specifically, we introduce a large, negative constant, $z$, to the unary choice payoffs $\pi_{\omega=\{i\}}(i)$ for all $i\in\mathcal{S}$. Let $\{\pi_{\omega=\{i\}}(i):i\in\mathcal{S}\}$ represent the set of unary choice payoffs, corresponding to the payoffs Alice receives when selecting a similarity class $i$ at a node $\omega$ where $\mathcal{S}_{\omega}=\{i\}$. We adjust each payoff in this set by adding the constant $z$, where $z$ is sufficiently negative, below a specified threshold $\tilde{z}<0$, i.e., $z<\tilde{z}<0$. In contrast, the payoffs $\{\pi_{\omega}(i):i\in\mathcal{S},\omega\in\Omega,|\mathcal{S}_{\omega}|>1\}$, associated with non-trivial choice nodes (where multiple similarity classes are available), remain unchanged.272727$\forall$ $i\in\mathcal{S}$, Alice is assumed to receive significantly lower payoffs when choosing a similarity class $i$ at a node where it is the sole available class, compared to nodes where at least one other similarity class is present.

### v2

@@ 3.1 Example: Multiplicity of SVE
    Figure 4: Stable Strict Pure SVE at $(1.0,0.0)$; $\beta=50$
    Figure 5: Stable Strict Pure SVE at $(0.0,0.5)$; $\beta=50$
@@ 3.3 Example: Unique Pure SVE
Figure 6: Stable Unique Mixed SVE at $(2.577,2.577)$; $\beta=50$
Figure 7: Stable Unique Strict Pure SVE at $(1.5,0)$; $\beta=50$
It is important to note that if the decision-maker possesses the finest partition available (i.e., she can distinguish each alternative independently), then multiple equilibria and mixed equilibria are impossible in a decision tree with generic payoffs in our learning model. Thus, the clustering of alternatives has non-trivial consequences on the dynamics of an otherwise standard reinforcement learning model. Finally, the examples suggest that an asymptotically stable equilibrium of the CPAL model always exists, and that strict pure equilibria are asymptotically stable whenever they are present. In the next section, we formalize these insights and provide general results on the convergence of CPAL dynamics when the DM operates with at most two similarity classes. Later in the paper, we analyze the general case with an arbitrary number of similarity classes. [^19] Supplementary examples showing the workings of the CPAL model with more than two similarity classes can be found in Sec. 3 of the Online Appendix.
@@ 4 Decision Trees with Two Similarity Classes
There exists a finite $\hat{\beta}$, such that $\forall$ $\beta>\hat{\beta}$, a decision tree $\mathcal{T^{\prime}}$ with generic payoffs where an agent chooses among alternatives in at most two similarity classes admits a smooth valuation equilibrium that is locally asymptotically stable under the CPAL dynamics. Additionally, if the equilibrium is unique, it is globally asymptotically stable.
@@ 5 Decision Trees with More than Two Similarity Classes
We extend our analysis to decision problems with generic payoffs where Alice has an arbitrary number of similarity classes available to her. Specifically, we let $|\mathcal{S}|=n$ where $n\in\mathbb{N}:n>2$, although we still maintain that $n$ is finite and that at any node $\omega$, $1\leq|\mathcal{S}_{\omega}|\leq n$. Once again, we imagine a decision tree where at the root $r^{\prime}$, nature selects a node $\omega\in\Omega$ with probability $p(\omega)$, where $\Omega={\mathcal{P(S)}}\setminus\{\emptyset\}$. For Theorem 3, we additionally assume that the set of nodes $\{\omega\in\Omega:|\mathcal{S}_{\omega}|=1\}$ is included in the support of the probability mass function $p$. That is, for each similarity class $s\in\mathcal{S}$, the trivial unary choice node featuring it is drawn by nature with a strictly positive probability. We recall that Theorem 1 guarantees the existence of a steady-state of the CPAL model and Lemma 2.1 tells us that a steady-state may arise in an arbitrarily small neighborhood of either a strict pure VE, a partially-mixed VE or a fully-mixed VE for a sufficiently large sensitivity parameter.
@@ 5.2 When Trivial Choice Payoffs are Large
The proof is relegated to Section 2 of the Online Appendix. ∎
@@ 5.3 When Trivial Choice Payoffs are Small
To illustrate the dramatic impact of assuming large trivial choice payoffs, we now consider the opposite scenario where these payoffs are small, leading to markedly different properties for the steady-states of the CPAL dynamics. Specifically, we introduce a large, negative constant, $z$, to the unary choice payoffs $\pi_{\omega=\{i\}}(i)$ for all $i\in\mathcal{S}$. Let $\{\pi_{\omega=\{i\}}(i):i\in\mathcal{S}\}$ represent the set of unary choice payoffs, corresponding to the payoffs Alice receives when selecting a similarity class $i$ at a node $\omega$ where $\mathcal{S}_{\omega}=\{i\}$. We adjust each payoff in this set by adding the constant $z$, where $z$ is sufficiently negative, below a specified threshold $\tilde{z}<0$, i.e., $z<\tilde{z}<0$. In contrast, the payoffs $\{\pi_{\omega}(i):i\in\mathcal{S},\omega\in\Omega,|\mathcal{S}_{\omega}|>1\}$, associated with non-trivial choice nodes (where multiple similarity classes are available), remain unchanged272727$\forall$ $i\in\mathcal{S}$, Alice is assumed to receive significantly lower payoffs when choosing a similarity class $i$ at a node where it is the sole available class, compared to nodes where at least one other similarity class is present..